OpenStudy (anonymous):

Look at the formula for finding distance. d = rt Which of the following is the formula to find time (t)?

OpenStudy (johnweldon1993):

well you have d = rt if you want to solve for t...you just rearrange the equation right? how would you isolate "t"?

OpenStudy (anonymous):

ooooo that makes sense now t = dr

OpenStudy (johnweldon1993):

not quite...try that again

OpenStudy (johnweldon1993):

remember that r is being MULTIPLIED to t...so to get t by itself you need to divide by r right?

OpenStudy (anonymous):

so t = d / r ? or t = r / d

OpenStudy (johnweldon1993):

well d = rt ------- r r d -- = t r Right?

OpenStudy (anonymous):

YUP :)

OpenStudy (johnweldon1993):

there ya go :)